OpenStudy (anonymous):

Solve the following equation. Then place the correct number in the box provided. |3- x | = 8 x = or x =

OpenStudy (anonymous):

Implicit in the way they ask you to put your answer, you know there will be two solutions. Do you know what the vertical bars ( | ) stand for?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

11 and -5 i think

OpenStudy (anonymous):

The vertical bars mean absolute value. When you're evaluating the expression, positive is positive and negative is positive. So, if I want to know |x|, then answer is 10 if x is 10 *or* -10.

OpenStudy (anonymous):

In |x| = A, you have x= A if x is positive. You have -x = A if x is negative. The second consideration is resolved by a double negative.

OpenStudy (anonymous):

o well i was wrong

OpenStudy (anonymous):

@mtmb11 No you weren't.

OpenStudy (anonymous):

So, if 3-x is positive, you have the equation 3 - x = 8 and then you solve that. If 3-x is negative, you have -(3 - x) = 8 or x - 3 = 8, and then you solve that one. You will have both as answers.

OpenStudy (anonymous):

yay i was correct

OpenStudy (anonymous):

how do i make my numbers go up on this account like right now i have 20

OpenStudy (anonymous):

So, do you understand now how you have to set up 2 equations? one for the positive and one for the negative.